### v1

# Mixed-state geometric phases of coherent and squeezed spin states

Abstract
Two mixed-state geometric phases, known as the Uhlmann phase and interferometric phase (IGP), of spin coherent states (CSSs) and spin squeezed states (SSSs) are analyzed. While each phase follows its parallel-transport condition, we also consider the non-adiabatic IGP for arbitrary unitary evolution beyond parallel transport. For the $j=3/2$ CSS, the Uhlmann phase shows temperature-induced topological phase transitions with jumps. The IGP and non-adiabatic IGP for the $j=3/2$ CSS also exhibits temperature-induced jumps. In contrast, the Uhlmann phase of the $j=1$ SSS exhibits smooth behavior without any temperature-induced transition. Interestingly, the parallel-transport condition of the IGP of the $j=1$ SSS in general does not allow a solution at finite temperature. Instead, the non-adiabatic IGP for the $j=1$ SSS has a solution showing smooth behavior as the squeezing parameter and temperature change. We also briefly discuss possible experimental implications and simulations.
I Introduction
Geometric phases have been a key concept in topological properties of quantum systems [1, 2]. The Berry phase arises when a quantum system undergoes adiabatic cyclic evolution in the context of pure quantum states and has been fundamental in studying topological insulators and superconductors, where the Berry curvature characterizes the underlying topological features [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. However, quantum systems are often in mixed states due to finite temperatures or non-equilibrium conditions, which call for generalizations of the Berry phase to mixed quantum states. A mathematically rigorous approach is the Uhlmann phase [17, 18, 19, 20] for mixed states. Like the Berry phase, the Uhlmann phase is obtained from parallel transport in a cyclic process, and the Uhlmann-Berry correspondence shows that the Uhlmann phase in the low-temperature limit generally agrees with the Berry phase of the ground state [21]. However, the Uhlmann bundle is trivial [22], which limits its applicability to defining other quantized topological indices. Nevertheless, the Uhlmann phase of bosonic and fermionic coherent states have been studied [21], in addition to that of two-level and spin-$j$ systems [23, 24, 25]. With a suitable generalization, the Uhlmann phase of dynamic systems has also been studied [26].
Another approach called the interferometric geometric phase (IGP) for mixed states has been proposed [27] and further analyzed [28, 29, 30, 31, 32], extending the geometric phase concept via an analogy of the optical Mach-Zehnder interferometer [33]. This IGP has been experimentally observed using nuclear magnetic resonance and polarized neutrons [34, 35, 36, 33]. Unlike the Uhlmann phase, the IGP does not require the evolution of the system to be cyclic [27]. In some cases, it can be shown that the IGP is the argument of the weighted sum of the Berry-phase factors of the constituent states [27].
Ref. [37] presents a comparison between the Uhlmann phase and IGP to contrast their differences. The Uhlmann phase, for instance, requires the ancillary system modeling environmental effects to be manipulated according to the change of the system and can exhibit finite-temperature topological phase transitions with quantized jumps in two-level systems. In contrast, the IGP does not exhibit such transitions at finite temperatures for general two-level systems. Recent studies have demonstrated that the IGP remains intact as temperature varies for the Kitaev chain [38], in contrast to the discrete jumps seen in the Uhlmann phase [23]. Nevertheless, quantized jumps of the IGP as temperature increases can occur in three-level or higher-dimensional systems [37]. Interestingly, these generalizations of the mixed-state geometric phases provide further insights into topological quantum phase transitions (TQPTs) and dynamical quantum phase transitions (DQPTs) [39].
Meanwhile, spin systems have served as paradigms in various physics branches, ranging from quantum mechanics [40] to statistical physics [41] to spintronic devices [42]. Two types of spin states have received particular attention: The coherent spin states (CSSs) are minimal uncertainty states of collective angular momentum. The CSS is the building block for the path-integral description of many-body spin systems [43] and serves as a classical approximation of quantum spin systems, thereby establishing the standard quantum limit (SQL) for measurements [44, 45, 46, 47]. However, surpassing the SQL is critical for advancing quantum metrology and quantum information technologies. Spin squeezing [48, 49, 50] allows for a reduction of the variance of one spin component at the expense of increasing that of another orthogonal direction, resulting in the squeezed spin states (SSSs). The SSSs enable precision measurements beyond the SQL and approach the Heisenberg limit with potential applications in, for example, atomic clocks [51, 52]. Significant improvements in precision measurement using spin squeezing have also been made in atomic Bose-Einstein condensates (BECs) [53, 54]. These developments not only highlight the metrological advantage of the SSS but also open possible avenues for generating quantum entanglement in many-body systems, fostering further advancement in quantum information processing [55].
Complementing the progress in surpassing the SQL through spin squeezing, here we set to investigate the geometric phases of the CSS and SSS at finite temperatures and search for topological phases or transitions. The possibility of finding different mixed-state geometric phases according to different evolution conditions already demonstrates rich physics in the underlying topology of finite-temperature systems. While the geometric phase of the ground-state CSS has been analyzed [56, 57], here we explicitly calculate the Uhlmann phase and the IGP for selected CSS and SSS at finite temperatures. We found that for the $j=3/2$ CSS, both Uhlmann phase and IGP demonstrate temperature-dependent topological or geometric phase transitions with quantized jumps. The jumps survive after generalizing the IGP to the non-adiabatic IGP for the CSS. In contrast, the $j=1$ SSS has a smoothly-varying Uhlmann phase as temperature increases, so there is no topological phase transition. More interestingly, the parallel-transport condition of the $j=1$ SSS cannot be satisfied at finite temperatures in general, leaving a construction of the IGP impractical. We therefore analyze the non-adiabatic IGP for the $j=1$ SSS to circumvent the obstacle but found no discrete jump in the geometric contribution of the phase at finite temperature.
The rest of the paper is organized as follows. Sec. II provides a self-contained theoretical background, covering the CSS, SSS, Uhlmann phase, and the adiabatic and non-adiabatic IGPs. Sec. III and Sec. IV present our systematic investigations of the mixed-state geometric phases for selected cases of the CSS and SSS, respectively. Section V discusses possible experimental and theoretical implications of our findings. Finally, Sec. VI concludes our work.
II THEORETICAL BACKGROUND
We begin with a brief overview of the CSS and SSS as well as a short review of the Uhlmann phase and interferometric geometric phase (IGP) of mixed states. Additionally, we will generalize the concept of IGP to allow arbitrary unitary evolution. For simplicity, we set $c=\hbar=k_{\text{B}}=1$ throughout the paper.
II.1 Overview of CSS
The spin operators satisfy the algebra [40]
$[J_{z},J_{+}]=J_{+},\quad[J_{z},J_{-}]=-J_{-},\quad[J_{+},J_{-}]=2J_{z}.$ (1)
The eigenstates of $J^{2}$ and $J_{z}$ are $|jm\rangle,\leavevmode\nobreak\ -j\leq m\leq j$. Moreover, $J_{\pm}|jm\rangle=\sqrt{(j\mp m)(j\pm m+1)}|jm\pm 1\rangle$ and $J_{z}|jm\rangle=m|jm\rangle$. The CSS is defined as [58, 59]
$|\xi\rangle:=\mathrm{e}^{\xi J_{+}-\bar{\xi}J_{-}}|j,-j\rangle,$ (2)
where $\xi\in\mathds{C}$ is a complex number.
Using the Baker-Campbell-Hausdorff disentangling formula [58], the CSS can also be expressed as
$|\xi\rangle$ $=\mathrm{e}^{\zeta J_{+}}\mathrm{e}^{\ln\left(1+|\zeta|^{2}\right)J_{z}}\mathrm{e}^{-\bar{\zeta}J_{-}}|j,-j\rangle$
$=\frac{1}{(1+|\zeta|^{2})^{j}}\mathrm{e}^{\zeta J_{+}}|j,-j\rangle\equiv|\zeta\rangle$ (3)
with $\zeta=\zeta(\xi)=\frac{\xi\tan|\xi|}{|\xi|}$. After defining $D(\xi)\equiv D(\zeta)=\mathrm{e}^{\xi J_{+}-\bar{\xi}J_{-}}$, the CSS corresponds to the ground state of the translated Hamiltonian $\hat{H}(\xi)\equiv\hat{H}(\zeta)=D(\zeta)\hat{H}D^{\dagger}(\zeta)$. Here $\hat{H}=\omega_{0}J_{z}$ , and $\omega_{0}$ is the Larmor frequency. The associated excited states are obtained in a similar manner: $|jm,\xi\rangle=|jm,\zeta\rangle=D(\zeta)|jm\rangle$, $-j\leq m\leq j$. By parameterizing $\xi$ as $\xi=\mathrm{e}^{-\mathrm{i}\phi}\frac{\theta}{2}$, we have $\zeta=\mathrm{e}^{-\mathrm{i}\phi}\tan\frac{\theta}{2}$. If the parameter ranges are chosen as $0\leq\phi\leq 2\pi$, $0\leq\theta\leq\pi$, then $\zeta$ spans the entire complex plane. Consequently, we will use $\zeta$ as the control parameter. The map $(\theta,\phi)\rightarrow\zeta$ projects the unit sphere onto the entire complex plane. Actually, the CSSs are closely related to the eigenstates of the corresponding spin-$j$ system with the Hamiltonian $H=\omega_{0}\hat{\mathbf{B}}\cdot\mathbf{J}$, where $\hat{\mathbf{B}}$ is the unit vector pointing along the direction of an external magnetic field [43].
II.2 Overview of SSS
[…]
$J_{\alpha}=\frac{1}{2}\sum_{l=1}^{N}\sigma_{l\alpha},\quad\alpha=x,y,z,$ (4)
where $\sigma_{l\alpha}$ is the Pauli matrix for the $l$-th object. Here we consider the spin-squeezed states (SSS) generated by the one-axis twisting Hamiltonian [49, 60]
$H_{\mathrm{OAT}}=\eta J_{x}^{2}=\frac{\eta}{4}\sum_{k,l=1}^{N}\sigma_{kx}\sigma_{lx},$ (5)
which is a nonlinear operator with the coupling constant $\eta$ involving all pairwise interactions. This indicates the presence of pairwise correlations in the spin-squeezed states generated by the Hamiltonian. Choosing the initial state as $|j,-j\rangle=|1\rangle^{\otimes N}$, where $|1\rangle=|j=1/2,j_{z}=1/2\rangle$, the SSS at time $t$ is formally written as:
$|\Psi(t)\rangle=\exp\left(-\mathrm{i}\Theta J_{x}^{2}/2\right)|1\rangle^{\otimes N},$ (6)
where $\Theta=2\eta t$.
We can generalize this concept of spin-squeezed states by acting the operator $\exp\left(-\mathrm{i}\Theta J_{x}^{2}/2\right)$ on $|jm\rangle$:
$|jm,\Theta\rangle=\exp\left(-\mathrm{i}\Theta J_{x}^{2}/2\right)|jm\rangle.$ (7)
By defining
$S(\Theta)=\mathrm{e}^{-\mathrm{i}\Theta J_{x}^{2}/2},$ (8)
then $|jm,\Theta\rangle$ is an eigenstate of the Hamiltonian
$H(\Theta)=S(\Theta)\omega_{0}J_{z}S^{\dagger}(\Theta).$ (9)
II.3 Uhlmann phase
A mathematically rigorous approach for generalizing the Berry phase from pure states to mixed states is the Uhlmann phase [17] through purification of density matrices, which can be described by the underlying fiber-bundle structure [61], An operator $W$ is said to purify the density matrix $\rho$ if $\rho=WW^{\dagger}$. Conversely, the purification $W$ can be uniquely decomposed as $W=\sqrt{\rho}\mathcal{U}$ if $\rho$ is full-rank (i.e., rank$(\rho)=N$ for a system with a $N$-dimensional Hilbert space). Here the unitary operator $\mathcal{U}$ is referred to as the phase factor of mixed states, representing a generalization of the U(1) phase factor associated with pure states.
We consider a system depending on a set of parameters $\mathbf{R}=(R_{1},R_{2},\cdots,R_{k})^{T}\in M$, where $M$ is the parameter manifold. When $\rho$ evolves along a smooth curve $C(t):=\mathbf{R}(t)$ in $M$, the purification of the density matrix also continuously evolves along an induced curve $\gamma(t):=W(t)=\sqrt{\rho(\mathbf{R}(t))}\mathcal{U}(\mathbf{R}(t))$. If the length of $\gamma$, defined as $L(\gamma)={\int}_{\gamma}\sqrt{\text{Tr}(\dot{W}\dot{W}^{\dagger})}\mathrm{d}t$, is minimized, $W$ is said to undergo a parallel-transport process. Therefore, $\gamma(t)$ is the horizontal lift of $\rho(t)\equiv\rho(\mathbf{R}(t))$. It can be shown that the Uhlmann parallel-transport condition is given by [62]
$\dot{W}W^{\dagger}=W\dot{W}^{\dagger}.$ (10)
This condition maximally ensures the “parallelity” between purification of adjacent states.
When $\rho$ experiences a cyclic evolution over a duration $\tau$, such that $\rho(0)=\rho(\tau)$, the evolution of its purification may not necessarily return to its initial value. Specifically, the initial and final phase factors of a Uhlmann process are related by
$\mathcal{U}(\tau)=\mathcal{P}\mathrm{e}^{-\oint_{C}A_{U}}\mathcal{U}(0),$ (11)
where $\mathcal{P}$ is the path-ordering operator and
$A_{U}=-\sum_{mn}|m\rangle\frac{\langle m|[\mathrm{d}\sqrt{\rho},\sqrt{\rho}]|n\rangle}{\lambda_{m}+\lambda_{n}}\langle n|$ (12)
is the Uhlmann connection. Here $\lambda_{n}$, $|n\rangle$ represent the eigenvalues and eigenstates of $\rho$, respectively. Notably, the parallelity lacks transitivity [39] even during parallel transport. Hence, the operator $\mathcal{P}\mathrm{e}^{-\oint_{C}A_{U}}$ associated with the Uhlmann holonomy quantifies a measure of the loss of parallelity. The following Uhlmann phase then reflects the topological change of the Uhlmann holonomy.
$\theta_{U}(C)=\arg\text{Tr}(W^{\dagger}(0)W(\tau))=\arg\text{Tr}[\rho(0)\mathcal{P}\mathrm{e}^{-\oint_{C}A_{U}}].$ (13)
II.4 Interferometric geometric phase
There is another approach for constructing a geometric phase for mixed states. Inspired by optical processes in the Mach-Zehnder interferometer, Ref. [27] assigned a phase to a mixed quantum state represented by the density matrix $\rho$ undergoing a unitary evolution $U(t)$. Explicitly, if a system evolves according to $\rho(t)=U(t)\rho(0)U^{\dagger}(t)$, it gives rise to a total phase
$\theta_{\text{T}}(t)=\arg\text{Tr}\left[\rho(0)U(t)\right].$ (14)
Notably, the transformation from $\rho(t)$ to $\rho(t+\mathrm{d}t)$ is $\rho(t+\mathrm{d}t)=U(t+\mathrm{d}t)\rho(0)U^{\dagger}(t+\mathrm{d}t)=U(t+\mathrm{d}t)U^{\dagger}(t)\rho(t)U(t)U^{\dagger}(t+\mathrm{d}t)$. Thus, by definition, the relative phase between the mixed states at $t$ and $t+\mathrm{d}t$ is $\theta_{\text{T}}(t)=\arg\text{Tr}\left[\rho(t)U(t+\mathrm{d}t)U^{\dagger}(t)\right]$. In this framework [27], a parallel-transport condition of $U(t)$ is proposed as follows. $\rho(t+\mathrm{d}t)$ is required to be instantaneously “in phase” with $\rho(t)$, which means $\arg\text{Tr}\left[\rho(t)U(t+\mathrm{d}t)U^{\dagger}(t)\right]=0$. In the differential form, the condition yields
$\text{Tr}\left[\rho(t)\dot{U}(t)U^{\dagger}(t)\right]=\text{Tr}\left[\rho(0)U^{\dagger}(t)\dot{U}(t)\right]=0.$ (15)
To specify parallel-transport by the evolution according to $U(t)$ more accurately, the condition (15) has been strengthened as [27]
$\langle n(t)|\dot{U}(t)U^{\dagger}(t)|n(t)\rangle=0,\quad n=0,1,2,\cdots,$ (16)
where $|n(t)\rangle$ is the $n$th eigen-vector of $\rho(t)$. We emphasize that the above parallel-transport condition is distinct from the Uhlmann parallel-transport condition (10) which depends explicitly on the choice of the U$(N)$ phase factor $\mathcal{U}$.
Under the parallel-transport condition, the total phase accumulated during the evolution is the geometric phase referred to as the IGP:
$\theta_{\text{G}}(t)=\theta_{\text{T}}(t)=\arg\text{Tr}\left[\rho(0)U(t)\right].$ (17)
This is because there is no accumulation of the dynamic phase during parallel transport, which can be understood by noting that if $U(t)$ represents a dynamic evolution process according to $\mathrm{i}\dot{U}=HU$, the parallel-transport condition (15) guarantees
$\theta_{\text{D}}(t)$ $=-\int_{0}^{t}\mathrm{d}t^{\prime}\text{Tr}\left[\rho(t^{\prime})H(t^{\prime})\right]$
$=-\mathrm{i}\int_{0}^{t}\mathrm{d}t^{\prime}\text{Tr}\left[\rho(t^{\prime})\dot{U}(t^{\prime})U^{\dagger}(t^{\prime})\right]=0.$ (18)
As a consequence, the total phase contains only the geometric contribution from the IGP.
In certain situations, it may be difficult for unitary evolutions to fully satisfy the parallel-transport condition (15). Nevertheless, the above discussion can be extended to general unitary evolution beyond parallel transport, which results in non-adiabatic phases [63]. For pure states, a similar formalism is known as the quantum kinematic approach to the geometric phase [64]. Recently, a similar non-adiabatic geometric phase for mixed states has been discussed in Ref. [65]. We follow those ideas and define the non-adiabatic IGP for any unitary process by subtracting the non-adiabatic generalization of the dynamic phase from the total phase. Explicitly,
$\theta_{\text{G}}^{\prime}(t)=\theta_{\text{T}}(t)-\theta_{\text{NG}}(t).$ (19)
Here
$\theta_{\text{NG}}(t)=-\mathrm{i}\int_{0}^{t}\mathrm{d}t^{\prime}\text{Tr}\left[\rho(t^{\prime})\dot{U}(t^{\prime})U^{\dagger}(t^{\prime})\right]$ (20)
is considered as a non-geometric phase because if $U(t)$ represents a dynamic evolution, $\theta_{\text{NG}}(t)$ reduces to the dynamic phase. Moreover, if $U(t)$ satisfies the parallel-transport condition, $\theta_{\text{NG}}(t)=0$, and the total phase equals to the geometric phase given by the IGP.
II.5 Difference between the two geometrical phases
We emphasize that the two aforementioned geometric phases for mixed states are physically and mathematically distinct. Although the theory of IGP can also be formulated equivalently using a purification formalism of density matrices [27], any phase factor $\mathcal{U}$ is ruled out in the derivation. Specifically, only the trivial decomposition $W=\sqrt{\rho}$ is permissible for the IGP because otherwise the condition (15) would be violated.
In contrast, the generation of the Uhlmann phase inherently relies on the phase factor $\mathcal{U}$ according to the underlying fiber-bundle structure. All purifications of a given density matrix $\rho$ thus span a U$(N)$ fiber space. Consequently, the Uhlmann phase is naturally and directly connected to the topological properties of the U$(N)$ principal bundle, a feature absent in the IGP. Explicitly, the Uhlmann phase corresponds to the Uhlmann holonomy of the Uhlmann bundle of density matrices, similar to the Berry phase revealing the Berry holonomy of the fiber bundle of pure states [61]. Moreover, the Uhlmann phase requires a cyclic process but is guaranteed to be gauge invariant [27, 37] while the IGP and its generalization do not impose those constraints. In the following, we will compare the differences between the Uhlmann phase and IGP through specific examples of the CSS and SSS.
III Geometrical Phases of CSS
III.1 The Uhlmann phase
To evaluate the Uhlmann phase, we consider a loop in the complex $\zeta$ plain: $C(t)\equiv\zeta(t)$, $0\leq t\leq\tau$ with $\zeta(0)=\zeta(\tau)$. The continuous transformation $D(\zeta(t))$ of a CSS at temperature $T$ induces a corresponding loop $\gamma(t):=\rho(\zeta(t))$ in the manifold of density matrices, where
$\rho(\zeta)=\frac{1}{Z}\mathrm{e}^{-\beta\hat{H}(\zeta)}=\frac{1}{Z}D(\zeta)\mathrm{e}^{-\beta\hat{H}}D^{\dagger}(\zeta).$ (21)
Here $\beta=1/(k_{B}T)$. Since $D(\zeta)$ is unitary, the eigenvalues of the Hamiltonian remain invariant under $D(\zeta)$. As a result, the partition function is a constant: $Z=\sum_{m=-j}^{j}\mathrm{e}^{-m\beta\omega_{0}}=\frac{\sinh\left[\left(j+\frac{1}{2}\right)\beta\omega_{0}\right]}{\sinh\left(\frac{\beta\omega_{0}}{2}\right)}$.
According to Eq. (12), the Uhlmann connection of the CSS is given by
$A_{U}=-\sum_{n\neq m}\frac{(\sqrt{\lambda_{n}}-\sqrt{\lambda_{m}})^{2}}{\lambda_{n}+\lambda_{n}}|jn,\zeta\rangle\langle jn,\zeta|\mathrm{d}|jm,\zeta\rangle\langle jm,\zeta|$
$=-\sum_{n\neq m}\chi_{nm}D(\zeta)|jn\rangle\langle jn|D^{\dagger}(\zeta)\mathrm{d}D(\zeta)|jm\rangle\langle jm|D^{\dagger}(\zeta),$ (22)
where $\lambda_{m}=\frac{1}{Z}\mathrm{e}^{-m\beta\omega_{0}}$ is the $m$th eigenvalue of the density matrix, and $\chi_{nm}=\frac{(\mathrm{e}^{-\frac{n}{2}\beta\omega_{0}}-\mathrm{e}^{-\frac{m}{2}\beta\omega_{0}})^{2}}{\mathrm{e}^{-n\beta\omega_{0}}+\mathrm{e}^{-m\beta\omega_{0}}}$. By applying Eq. (II.1), we further get
$D^{\dagger}(\zeta)\mathrm{d}D(\zeta)=\frac{J_{+}\mathrm{d}\zeta-J_{-}\mathrm{d}\bar{\zeta}}{1+|\zeta|^{2}}+J_{z}\frac{\zeta\mathrm{d}\bar{\zeta}-\bar{\zeta}\mathrm{d}\zeta}{1+|\zeta|^{2}}.$ (23)
Accordingly, the Uhlmann connection becoms
$A_{U}=-\sum_{n\neq m}\frac{\chi_{nm}}{1+|\zeta|^{2}}\times$
$D(\zeta)|jn\rangle\langle jn|\left(J_{+}\mathrm{d}\zeta-J_{-}\mathrm{d}\bar{\zeta}\right)|jm\rangle\langle jm|D^{\dagger}(\zeta)$
$=-\frac{\chi}{(1+|\zeta|^{2})^{2}}\left[J_{+}(\mathrm{d}\zeta+\zeta^{2}\mathrm{d}\bar{\zeta})-J_{-}(\mathrm{d}\bar{\zeta}+\bar{\zeta}^{2}\mathrm{d}\zeta)\right.$
$\left.+2J_{z}(\zeta\mathrm{d}\bar{\zeta}-\bar{\zeta}\mathrm{d}\zeta)\right],$ (24)
where $\chi\equiv\chi_{n+1,n}=\chi_{n-1,n}=1-\text{\text{sech}}\frac{\beta\omega_{0}}{2}$. The $n=m$ terms have been included to simplify the expression since $\chi_{nn}=0$ gives no contributions.
Next, noting that $\mathrm{d}\zeta=-\mathrm{i}\mathrm{d}\phi\zeta+\frac{1}{2}\mathrm{d}\theta\mathrm{e}^{-\mathrm{i}\phi}\sec^{2}\frac{\theta}{2}$, we obtain the final expression of $A_{U}$:
$A_{U}$ $=\mathrm{i}\chi\left[\left(J_{x}\cos\phi+J_{y}\sin\phi\right)\cos\theta-J_{z}\sin\theta\right]\sin\theta\mathrm{d}\phi$
$\quad+\mathrm{i}\chi\left(J_{x}\sin\phi-J_{y}\cos\phi\right)\mathrm{d}\theta.$ (25)
Using Eq. (III.1), the Uhlmann phase generated in a cyclic process can be evaluated. The $A_{U}$ of the CSS only differs by a minus sign from that of the spin-$j$ system discussed in Ref. [66]. Therefore, their Uhlmann phases only differ by a minus sign. As explained in Sec. II, this is because the CSS is related to the spin-$j$ system with a particular Hamiltonian. Since the cases with $j=\frac{1}{2}$ and $1$ of the corresponding spin-$j$ systems have already been studied in Ref. [66], we will focus here on the $j=\frac{3}{2}$ CSS. In Appendix A, we show that (I) the Uhlmann phase reduces to the Berry phase associated with the ground state as $T\rightarrow 0$, thereby confirming the correspondence explained in Ref. [21], and (II) the $T\rightarrow 0$ expressions of the CSSs agree with those of Ref. [56] for pure states.
To explicitly show the properties of the Uhlmann phase of a $j=\frac{3}{2}$ CSS, we choose the case evolving along the equator ($\theta=\frac{\pi}{2}$) in the parameter space. In this case, the Uhlmann connection reduces to $A_{U}=-\mathrm{i}\chi J_{z}\mathrm{d}\phi$. The spin operators are
$J_{x}$ $=\left(\begin{array}[]{cccc}0&\frac{\sqrt{3}}{2}&0&0\\ \frac{\sqrt{3}}{2}&0&1&0\\ 0&1&0&\frac{\sqrt{3}}{2}\\ 0&0&\frac{\sqrt{3}}{2}&0\\ \end{array}\right),$ (30)
$J_{y}$ $=\left(\begin{array}[]{cccc}0&\frac{\mathrm{i}\sqrt{3}}{2}&0&0\\ -\frac{\mathrm{i}\sqrt{3}}{2}&0&\mathrm{i}&0\\ 0&-\mathrm{i}&0&\frac{\mathrm{i}\sqrt{3}}{2}\\ 0&0&-\frac{\mathrm{i}\sqrt{3}}{2}&0\\ \end{array}\right),$ (35)
$J_{z}$ $=\left(\begin{array}[]{cccc}-\frac{3}{2}&0&0&0\\ 0&-\frac{1}{2}&0&0\\ 0&0&\frac{1}{2}&0\\ 0&0&0&\frac{3}{2}\\ \end{array}\right).$ (40)
We also assume that the evolution starts from $\phi=0$, so the initial density matrix is given by $\rho(0)=\frac{1}{Z}\mathrm{e}^{\frac{\mathrm{i}\pi}{2}J_{y}}\mathrm{e}^{-\beta\omega_{0}J_{z}}\mathrm{e}^{-\frac{\mathrm{i}\pi}{2}J_{y}}$ according to Eq. (21). A straightforward evaluation shows that the Uhlmann phase generated in this process is
$\theta_{U}$ $=\arg\operatorname{Tr}\left[\frac{1}{Z}\mathrm{e}^{\beta\omega_{0}J_{x}}\mathcal{P}e^{\mathrm{i}\chi{\oint}_{\gamma}J_{z}\mathrm{d}\phi}\right],$ (41)
where we have used the fact $\mathrm{e}^{\mathrm{i}\theta J_{y}}J_{z}\mathrm{e}^{-\mathrm{i}\theta J_{y}}=-J_{x}\sin\theta+J_{z}\cos\theta$.
Figure 1: Uhlmann phase $\theta_{U}$ versus temperature for the CSS with $j=\frac{3}{2}$ along $\theta=\frac{\pi}{2}$.
We plot $\theta_{U}$ as a function of $T$ in Figure 1. In the low temperature limit, $\theta_{U}=\pi$, which indicates that the topological properties of the system change after a cyclic evolution along the equator. More specifically, the final purification (in terms of the fiber in the bundle language) is “antiparallel” to the initial purification. As temperature increases and crosses the first critical temperature $T^{1}_{c}\approx 0.321\omega_{0}$, the value of $\theta_{U}$ suddenly drops to 0, implying that a topological change occurs in the evolution along the equator on the parameter space. The final purification now becomes “parallel” to the initial purification. As temperature continues to rise, subsequent topological transitions occur at $T^{2}_{c}\approx 0.376\omega_{0}$ and $T^{3}_{c}\approx 0.493\omega_{0}$. In the high temperature limit, $\theta_{U}$ remains zero, meaning the associated Uhlmann holonomy is topologically trivial. This is reasonable since $\lim_{T\rightarrow 0}\rho=\frac{1}{4}1_{4\times 4}$. Therefore, the evolution loop $\rho(t)$ essentially becomes a single point with its horizontal lift $W(t)$ also reduced to a single point [66]. Hence, there is no difference between the initial and final fibers anymore. We mention that bosonic and fermionic coherent states at finite temperatures, in contrast, have been shown to have smooth behavior of the Uhlmann phase without any jumps [21].
III.2 The interferometric geometric phase
Here we first evaluate the IGP for the CSS following unitary evolution that satisfies the parallel-transport condition (15). A natural candidate is the transformation that generates the CSS, as shown in Eq. (21). This yields
$\rho(t)\equiv\rho(\zeta(t))=D(\zeta(t))\rho(0)D^{\dagger}(\zeta(t)),$ (42)
where $\rho(0)=\frac{1}{Z}\mathrm{e}^{-\beta\hat{H}}$. Initially, we set $\zeta(0)=0$, implying $\theta(0)=0$. Using Eq. (23), the parallel-transport condition (15) can be satisfied by requiring
$\operatorname{Tr}\left[\rho(0)D^{\dagger}(\zeta)\mathrm{d}D(\zeta)\right]$
$=$ $\sum_{m}\frac{\mathrm{e}^{-m\beta\omega_{0}}}{Z}\left\langle jm\right|\frac{J_{+}\mathrm{d}\zeta-J_{-}\mathrm{d}\bar{\zeta}+J_{z}(\zeta\mathrm{d}\bar{\zeta}-\bar{\zeta}\mathrm{d}\zeta)}{1+|\zeta|^{2}}\left|jm\right\rangle$
$=$ $\frac{2\mathrm{i}}{Z(1+|\zeta|^{2})}\sum_{m=-j}^{j}m\mathrm{e}^{-m\beta\omega_{0}}\left(\tan^{2}\frac{\theta}{2}\mathrm{d}\phi\right)$
$=$ $0$ (43)
A possible choice is $\tan^{2}\frac{\theta}{2}\mathrm{d}\phi=0$, which implies either $\theta=0$ or $\mathrm{d}\phi=0$. The former leads to $\zeta=0$, indicating that $D(\zeta)=1$, which is a trivial evolution. Therefore, we choose the latter and set $\phi$ to be constant. This represents a longitude on the unit-sphere parameter space shown by the big circle in Figure 2.
Figure 2: The great circle represents a longitude with constant $\phi$ on the unit-sphere parameter space, which satisfies the parallel-transport condition (III.2). In contrast, the twisting curve denotes a simple evolutionary path with $\theta=\phi$, which does not fulfill the parallel-transport condition. Both paths start from the north pole, depicted by the top blue point.
To explicitly show the properties of the IGP along a path of fixed $\phi$, we set $j=\frac{3}{2}$ as the case of the Uhlmann phase and found
$\theta_{\text{G}}(t)$ $=\arg\operatorname{Tr}[\rho(0)D(t)]$
$=\arg\operatorname{Tr}[\frac{1}{Z}\mathrm{e}^{-\beta\omega_{0}J_{z}}\mathrm{e}^{\zeta(t)J_{+}}\mathrm{e}^{\ln\left(1+|\zeta(t)|^{2}\right)J_{z}}\mathrm{e}^{-\bar{\zeta}(t)J_{-}}]$
$=\arg\left[-2e^{\beta\omega_{0}}\tan^{2}\left(\frac{\theta(t)}{2}\right)+e^{2\beta\omega_{0}}+1\right].$ (44)
Interestingly, the result is independent of $\phi$. To visualize the behavior of the IGP, we introduce $\theta_{f}\equiv\theta(t_{\text{final}})$ as the final value of $\theta$ at the end of evolution by $D(t)$. In the top panel of Figure 3, we present the contour plot of the IGP as a function of $T$ and $\theta_{f}$ for $j=\frac{3}{2}$. One observes that when $\theta_{f}\in(\frac{\pi}{2},\frac{3\pi}{2})$, the value of the IGP undergoes a sudden jump from 0 to $\pi$ as the system crosses a critical temperature $T_{c}$. This means that $\rho(t_{\text{final}},T_{c}^{-})$ is in phase with $\rho(0,T_{c}^{-})$, whereas $\rho(t_{\text{final}},T_{c}^{+})$ is out of phase with $\rho(0,T_{c}^{+})$.
In the bottom panel of Figure 3, we choose a special case with $\theta_{f}=\frac{3}{4}\pi$ to show the discontinuity of the IGP at $T_{c}=0.408\omega_{0}$. Similar to the case of the Uhlmann phase, this can be recognized as a geometric phase transition induced by temperature. The only difference is that, in general, the IGP is not directly related to the topological properties of the system. Furthermore, finite-temperature geometric phase transitions of the IGP seem to be less common than those of the Uhlmann phase. For a two-level Hermitian quantum system, a finite change in temperature in general does not induce any jump of the IGP [67, 37]. For a three-level system, a careful design of the evolution process can make it happen [37]. Our results of the CSS thus illustrate additional transitions of the IGP and show that the CSS is suitable for studying interesting properties of mixed-state geometric phases.
Figure 3: (Top panel) Contour plot of the IGP $\theta_{G}$ versus temperature $T$ and $\theta_{f}$ for $j=\frac{3}{2}$. The geometric phase transitions when $\theta_{f}\in(\frac{\pi}{2},\frac{3\pi}{2})$ are indicated by the boundaries separating the two values of $\theta_{G}$. (Bottom panel) $\theta_{G}$ versus temperature $T$ for the case with $\theta_{f}=\frac{3\pi}{4}$.
The above studies can also be extended to generic unitary evolutions which do not necessarily satisfy the parallel-transport condition (III.2). As previously noted, the non-adiabatic IGP is generated during these evolutions according to Eq. (19):
$\theta^{\prime}_{\text{G}}(t)=\arg\operatorname{Tr}[\rho(0)D(t)]+\mathrm{i}\int_{0}^{t}\operatorname{Tr}\left[\rho(0)D^{\dagger}(t^{\prime})\dot{D}(t^{\prime})\right]\mathrm{d}t^{\prime}.$ (45)
The evolution is characterized by a trajectory $\theta=f(\phi)$ or, conversely, $\phi=g(\theta)$ in the parameter space. To contrast with the IGP, we focus on evolution during which $\theta^{\prime}_{\text{G}}$ exhibits finite jumps. Specifically, a simple evolutionary curve $C$ defined by $\theta(t)=\phi(t)$ with $\phi\in[0,\pi]$ is selected, as shown by the twisting curve in Fig. 2. This curve starts at $\theta=0$ when $t=0$ and ends at any point $(\theta_{f},\phi_{f})$ along the path, where $\theta_{f}\equiv\theta(t_{\text{final}})$ and $\phi_{f}\equiv\phi(t_{\text{final}})$. In the top panel of Figure 4, we present the contour plot of $\theta^{\prime}_{\text{G}}$ as a function of $T$ and $\phi_{f}$ for the $j=\frac{3}{2}$ CSS. Although $\theta^{\prime}_{\text{G}}$ is not quantized, its value still undergoes quantized jumps. To better illustrate those jumps, we plot $\theta^{\prime}_{\text{G}}$ vs $T$ with $\theta_{f}=\frac{3\pi}{4}$ and $\theta^{\prime}_{\text{G}}$ vs $\theta_{f}$ with $T=0.6\omega_{0}$, respectively, in the bottom panels of Fig. 4.
Interestingly, the critical temperature $T_{c}=0.408\omega_{0}$ with a $\pi$-jump when $\theta_{f}=\frac{3\pi}{4}$ is identical to that observed along the path with constant $\phi$ which respects the parallel-transport condition, as shown in the bottom panel of Fig. 3. This can be understood by noting that in the non-geometric phase, the second term on the right-hand-side of Eq. (45) is always continuous and vanishes during parallel transport. The discontinuity can only arise from the first term, which appears in both cases with and without the parallel-transport condition. At fixed temperature $T=0.6\omega_{0}$ but with varying end points labeled by $\theta_{f}$, $\theta^{\prime}_{\text{G}}$ also experiences a $\pi$-jump at about $\theta_{f}\approx 2.05$ rad, as indicated by the right panel in the bottom of Fig. 4. Therefore, the IGP and its non-adiabatic generalization of the CSS offer interesting predictions for future experimental investigations.
Figure 4: (Top panel) The non-adiabatic IGP $\theta^{\prime}_{\text{G}}$ as a function of $T$ and $\phi_{f}$, where $\theta_{f}=\phi_{f}$ corresponding to $\theta_{f}\in[0,\frac{3\pi}{4}]$. The black curve indicates geometrical phase transitions. (Bottom panels) $\theta^{\prime}_{\text{G}}$ vs $T$ at $\theta_{f}=\frac{3\pi}{4}$ (left) and $\theta^{\prime}_{\text{G}}$ vs $\theta_{f}$ at $T=0.6\omega_{0}$ (right). In both two cases, $\theta^{\prime}_{\text{G}}$ exhibits $\pi$-jumps.
IV GEOMETRICAL PHASES OF SSS
IV.1 Uhlmann Phase
For the SSS shown in Eq. (5), the parameter space is one-dimensional. Therefore, we need to determine the value of $\Theta$ to define a cyclic process for density matrices, which is essential for generating the Uhlmann phase. The density matrix of a system in thermal equilibrium governed by $H(\Theta)$ show in Eq. (9) is given by $\rho(\Theta)=\frac{1}{Z}\mathrm{e}^{-\beta H(\Theta)}$ with $Z=\operatorname{Tr}(\mathrm{e}^{-\beta H(\Theta)})$. By examining Eq. (6), we observe that the quadratic term of $J_{x}$ appears in the exponent, making it difficult to derive a general expression for the SSS. Therefore, for a concise illustration of the geometric phases of the SSS, we will focus our studies on a fixed value of $j$. Since the behavior of the IGP for a two-level system ($j=\frac{1}{2}$) is relatively simple [67], we will begin with $j=1$ for our analysis.
For the particular case of the $j=1$ SSS, the corresponding angular momentum operators are
$J_{x}$ $=\left(\begin{array}[]{ccc}0&\frac{1}{\sqrt{2}}&0\\ \frac{1}{\sqrt{2}}&0&\frac{1}{\sqrt{2}}\\ 0&\frac{1}{\sqrt{2}}&0\\ \end{array}\right),J_{y}=\left(\begin{array}[]{ccc}0&\frac{\mathrm{i}}{\sqrt{2}}&0\\ -\frac{\mathrm{i}}{\sqrt{2}}&0&\frac{\mathrm{i}}{\sqrt{2}}\\ 0&-\frac{\mathrm{i}}{\sqrt{2}}&0\\ \end{array}\right),$ (52)
$J_{z}$ $=\left(\begin{array}[]{ccc}-1&0&0\\ 0&0&0\\ 0&0&1\\ \end{array}\right).$ (56)
From Eq. (8), we have
$S(\Theta)=\left(\begin{array}[]{ccc}\frac{1}{2}+\frac{1}{2}e^{-\frac{\mathrm{i}\Theta}{2}}&0&-\frac{1}{2}+\frac{1}{2}e^{-\frac{\mathrm{i}\Theta}{2}}\\ 0&e^{-\frac{\mathrm{i}\Theta}{2}}&0\\ -\frac{1}{2}+\frac{1}{2}e^{-\frac{\mathrm{i}\Theta}{2}}&0&\frac{1}{2}+\frac{1}{2}e^{-\frac{\mathrm{i}\Theta}{2}}\\ \end{array}\right).$ (60)
A straightforward calculation shows that when $\Theta=4\pi$, $S(4\pi)=\mathbf{1}_{3\times 3}$, which leads to $\rho(4\pi)=\rho(0)$ and gives a cyclic process. The $4\pi$ periodicity also contrasts the $j=1$ SSS from an ordinary spin-1 state.
According to Eq. (12), the Uhlmann connection of the $j=1$ SSS is
$A_{U}$ $=-\sum_{n\neq m}\chi_{nm}S\left|jn\right\rangle\left\langle jn\right|S^{\dagger}\mathrm{d}S\left|jm\right\rangle\left\langle jm\right|S^{\dagger}$
$=\sum_{n\neq m}\chi S\left|jn\right\rangle\left\langle jn\right|\frac{\mathrm{i}\left(J^{2}_{+}+J^{2}_{-}\right)}{8}\left|jm\right\rangle\left\langle jm\right|S^{\dagger}\mathrm{d}\Theta$
$=\frac{\mathrm{i}\chi}{4}\left[J^{2}_{x}-SJ^{2}_{y}S^{\dagger}\right]\mathrm{d}\Theta,$ (61)
where $\chi_{nm}=\frac{\mathrm{e}^{-\beta m\omega_{0}}+\mathrm{e}^{-\beta n\omega_{0}}-2\mathrm{e}^{-\frac{\beta(m+n)\omega_{0}}{2}}}{\mathrm{e}^{-\beta m\omega_{0}}+\mathrm{e}^{-\beta n\omega_{0}}}$, and $\chi\equiv\chi_{n,n+2}=\chi_{n+2,n}=\frac{(\mathrm{e}^{\frac{\beta\omega_{0}}{2}}-\mathrm{e}^{\frac{-\beta\omega_{0}}{2}})^{2}}{\mathrm{e}^{\beta\omega_{0}}+\mathrm{e}^{-\beta\omega_{0}}}$. In the last line, we have included the vanishing $n=m$ and $n=m-1$ terms. It is in general challenging to find a nontrivial evolution path along which $A_{U}$ is proportional to a constant matrix. The difficulty thus hinders the discovery of an analytical expression of the Uhlmann phase.
Figure 5: Uhlmann phase $\theta_{U}$ as a function of temperature $T$ on semi-log scale. The smooth curve indicates the absence of a transition in this case.
Nevertheless, numerical results can be obtained by using the Trotter-Suzuki approximation [68, 69] to tackle the path-ordered expression. Our findings are presented in Figure 5, where $\theta_{U}$ is plotted as a function of $T$ on semi-log scale. Unlike the Uhlmann phase of the CSS, the Uhlmann phase of the SSS exhibits no discontinuities as the temperature varies, indicating the absence of a temperature-induced topological phase transition. The results show that at finite temperature, the initial and final purifications no longer remain parallel, but the “angle” between them changes continuously with temperature without any abrupt transition from “parallel” to “antiparallel” as seen previously in the CSS. In the infinite temperature limit, $\theta_{U}=2\pi\equiv 0\mod 2\pi$ again. This is reasonable since $\lim_{T\rightarrow+\infty}\rho$ is proportional to the identity matrix, and the associated horizontal lift shrinks to a single point, resulting in $\lim_{T\rightarrow+\infty}\theta_{U}=0\mod 2\pi$ [24].
IV.2 The interferometric geometric phase
To study the IGP of the SSS, we consider the unitary evolution
$\rho(\Theta(t))=\frac{1}{Z}\mathrm{e}^{-\beta H(\Theta(t))}=S(\Theta(t))\rho(\Theta(0))S^{\dagger}(\Theta(t)),$ (62)
where $\rho(\Theta(0))=\rho(0)=\frac{1}{Z}\mathrm{e}^{-\beta H}$. To ensure that $S$ satisfies the parallel-transport condition (15), it is required that
00 0 $=\operatorname{Tr}\left[\rho(0)S^{\dagger}\dot{S}\right]$
$=\sum_{m=-j}^{j}\left\langle jm\right|\frac{\mathrm{e}^{-\beta\omega_{0}J_{z}}}{Z}\frac{2J_{-}J_{+}+2J_{z}}{8}\left|jm\right\rangle$
$=\sum_{m=-j}^{j}\frac{\mathrm{e}^{-m\beta\omega_{0}}}{Z}\frac{j^{2}+j-m^{2}}{4}.$ (63)
However, this condition cannot be satisfied at finite temperatures in general, so the IGP in its original form is not guaranteed. Therefore, the SSS provides an explicit example where a compatible evolution for generating the IGP is missing. This is in contrast to the Uhlmann phase based on the underlying Uhlmann bundle, which makes it possible to find the parallel transport on the geometric structure through the construction of the Uhlmann connection.
[…]
Nevertheless, one can still obtain the non-adiabatic IGP based on Eq. (19) accumulated during an arbitrary evolution beyond the parallel-transport condition (15). To fairly compare with the results of the Uhlmann phase of the SSS, in the following we choose $j=1$ for the non-adiabatic IGP as well. The total phase of the $j=1$ SSS is given by
$\theta_{\text{T}}(t)=\arg\operatorname{Tr}[\rho(0)S(\Theta(t))]$
$=$ $\arg\left[\left(\frac{1}{2}+\frac{1}{2}\mathrm{e}^{-\frac{\mathrm{i}\Theta(t)}{2}}\right)\left(\mathrm{e}^{-\beta\omega_{0}}+\mathrm{e}^{\beta\omega_{0}}\right)+\mathrm{e}^{-\frac{\mathrm{i}\Theta(t)}{2}}\right],$ (64)
and the non-geometric phase is
$\theta_{\text{NG}}(t)$ $=-\mathrm{i}\int_{0}^{t}\operatorname{Tr}[\rho(0)S^{\dagger}(t^{\prime})\dot{S}(t^{\prime})]\mathrm{d}t^{\prime}$
$=-\frac{2+\mathrm{e}^{-\beta\omega_{0}}+\mathrm{e}^{\beta\omega_{0}}}{1+\mathrm{e}^{-\beta\omega_{0}}+\mathrm{e}^{\beta\omega_{0}}}\frac{\Theta(t)}{4}.$ (65)
The non-adiabatic IGP is then obtained by $\theta^{\prime}_{\text{G}}(t)=\theta_{\text{T}}(t)-\theta_{\text{NG}}(t)$. Since the evolution in this case is not necessarily periodic, $\theta_{\text{G}}$ also depends on the final (instantaneous) time $t$.
In the top panel of Figure 6, the non-adiabatic IGP $\theta_{\text{G}}^{\prime}$ is plotted as a function of temperature $T$ and the squeezing parameter $\Theta_{f}\equiv\Theta(t_{\text{final}})$. Similar to the Uhlmann phase of the $j=1$ SSS, there is no temperature-induced geometric phase transition here. We show the behavior of $\theta^{\prime}_{\text{G}}$ as a function of $T$ for $\Theta_{f}=5.4$ in the bottom panel, in which $\theta^{\prime}_{\text{G}}$ continuously decreases with $T$. At finite temperatures, we find that $\theta^{\prime}_{\text{G}}$ also continuously varies with $\Theta_{f}$ without any discontinuities.
However, in the zero-temperature limit, $\theta_{\text{T}}=\arg[1+\mathrm{e}^{-\frac{\mathrm{i}\Theta_{f}}{2}}]$ while $\theta_{\text{NG}}$ is always a continuous function of $\beta$ and $\Theta_{f}$. Therefore, as $\Theta_{f}$ crosses $2\pi$, the midpoint of the parameter range $[0,4\pi]$, the value of $\theta^{\prime}_{\text{G}}$ experiences a $\pi$-jump (from $\frac{\pi}{2}$ to $-\frac{\pi}{2}$ or vice versa), indicating a geometric phase transition. We show this behavior of $\theta^{\prime}_{\text{G}}$ in the middle-left panel of Figure 6. For comparison, we also plot $\theta^{\prime}_{\text{G}}$ vs $\Theta_{f}$ at $T=1.0\omega_{0}$ in the middle-right panel to contrast the smoothing of the curve at finite temperature. For the $j=1$ SSS, therefore, finite-temperature jumps of the two types of mixed-state geometric phases are absent.
V Implications
Since the CSS and SSS have been realized in condensed matter or atomic, molecular, and optical systems [53, 54, 70, 71, 72], experimental generation and measurement of the mixed-state geometric phases may become feasible in the future. However, imposing the parallel-transport condition for the IGP or Uhlmann phase requires precise controls of the systems and the ancilla representing the environment, which can be challenging in laboratories. Nevertheless, the non-adiabatic IGP relaxes the constraint and becomes a prominent candidates for demonstrating interesting geometric effects of CSS and SSS at finite temperatures. We caution that the non-adiabatic IGP is buried in the total phase, and a systematic removal of the non-geometric contribution according to Eq. (19) is required to obtain the genuine geometric contribution.
On the other hand, the rapid development of quantum computers opens possibilities of simulating complex quantum systems. By using an ancilla qubit acting as the environment and coupling it to a system qubit, Ref. [73] shows that the Uhlmann phase of a spin-$1/2$ system can be simulated on a quantum computer by using two engineered Hamiltonians to respectively evolve the system and ancilla simultaneously. Since a spin-$j$ object may be decomposed as a collection of $2j$ spin-$1/2$ objects, one may use $2j$ qubits as a system of a spin-$j$ system coupled to an ancilla system of $2j$ qubits to realize the CSS or SSS coupled to the environment. It then follows a series of technical challenges to construct suitable engineered Hamiltonians for the system and ancilla in order to generate the corresponding IGP or Uhlmann phase when the corresponding parallel-transport condition is imposed during the evolution. After the targeted geometric phase is generated on a quantum computer, one may follow Ref. [73] to couple additional measurement qubits to the composite system and perform state tomography to extract the phase accumulated during the constrained evolution.
VI Conclusion
We have derived the Uhlmann phase and the IGP along with its non-adiabatic generalization for the CSS and SSS. Selected examples are fully solved to demonstrate interesting properties of those mixed-state geometric phases. The Uhlmann phase is well-defined for both CSS and SSS due to its rigorous mathematical structure, but there may or many not be topological phase transitions manifested by discrete jumps of the Uhlmann phase as temperature varies. In contrast, the parallel-transport condition for the IGP does not have a general solution for the SSS although the IGP of the CSS exhibits discrete jumps. The non-adiabatic IGP forgoes the constraint and exhibits discrete jumps in the geometric part of the accumulated phase for the CSS. However, no finite-temperature jump is found in the non-adiabatic IGP for the $j=1$ SSS. Our formalisms provide the foundation for future theoretical and experimental investigations of mixed-state topological properties of quantum spin systems.
ACKNOWLEDGEMENTS
H.G. was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904) and the National Natural Science Foundation of China (Grant No. 12074064). X.Y.H. was supported by the Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No. 2023ZB611). C.C.C. was supported by the National Science Foundation under Grant No. PHY-2310656.
Appendix A Correspondence between Uhlmann phase and Berry phase of CSS as $T\rightarrow 0$
To evaluate the Uhlmann phase of the CSS, we introduce $g(t)=\mathcal{P}\mathrm{e}^{-{\int}_{0,\gamma}^{t}A_{U}(t^{\prime})}$, which is an operator depending on the evolution curve $\gamma$. Since $\gamma(\tau)=\gamma(0)$, $g(\tau)$ is in fact the Uhlmann holonomy. The operator $g(t)$ satisfies the differential equation
$\frac{\mathrm{d}g(t)}{\mathrm{d}t}=\frac{\chi}{1+|\zeta|^{2}}D(\zeta(t))(J_{+}\dot{\zeta}-J_{-}\dot{\bar{\zeta}})D^{\dagger}(\zeta(t))g(t)$ (66)
subject to the initial condition $g(0)=1$. To solve this equation, we define $g^{\prime}(t)=D^{\dagger}(\zeta(t))g(t)$. By taking the complex conjugation of both sides of Eq. (23), we get
$\dot{D}^{\dagger}D=\frac{1}{1+|\zeta|^{2}}\left[J_{-}\dot{\bar{\zeta}}-J_{+}\dot{\zeta}+J_{z}\left(\bar{\zeta}\dot{\zeta}-\zeta\dot{\bar{\zeta}}\right)\right].$ (67)
Using Eqs. (66) and (67), we have
$\frac{\mathrm{d}g^{\prime}(t)}{\mathrm{d}t}=\dot{D}^{\dagger}Dg^{\prime}(t)+\frac{\chi}{1+|\zeta|^{2}}\left(J_{+}\dot{\zeta}-J_{-}\dot{\bar{\zeta}}\right)g^{\prime}(t)$
$=\frac{1}{1+|\zeta|^{2}}\left[-\eta\left(J_{+}\dot{\zeta}-J_{-}\dot{\bar{\zeta}}\right)+J_{z}\left(\bar{\zeta}\dot{\zeta}-\zeta\dot{\bar{\zeta}}\right)\right]g^{\prime}(t),$
where $\eta=\text{\text{sech}}\frac{\beta\omega_{0}}{2}$. Solving this equation, we obtain
$g(\tau)=$ $D(\zeta(\tau))\mathcal{P}\mathrm{e}^{\frac{{\oint}_{\gamma}\left[-\eta\left(J_{+}\mathrm{d}\zeta-J_{-}\mathrm{d}\bar{\zeta}\right)+J_{z}\left(\bar{\zeta}\mathrm{d}\zeta-\zeta\mathrm{d}\bar{\zeta}\right)\right]}{1+|\zeta|^{2}}}D^{\dagger}(\zeta(0)).$ (68)
Therefore, the Uhlmann phase of the CSS is
$\theta_{U}$ $=\arg\text{Tr}\left[\rho(\zeta(0))g(\tau)\right]$
$=\arg\text{Tr}\left[\rho(0)\mathcal{P}\mathrm{e}^{\frac{{\oint}_{\gamma}\left[-\eta\left(J_{+}\mathrm{d}\zeta-J_{-}\mathrm{d}\bar{\zeta}\right)+J_{z}\left(\bar{\zeta}\mathrm{d}\zeta-\zeta\mathrm{d}\bar{\zeta}\right)\right]}{1+|\zeta|^{2}}}\right],$ (69)
where $\rho(0)=\frac{1}{Z}\mathrm{e}^{-\beta\omega_{0}J_{z}}$. As $T\rightarrow 0$, $\lim_{\beta\rightarrow\infty}\eta=\lim_{\beta\rightarrow\infty}\text{\text{sech}}\frac{\beta\omega_{0}}{2}=0$, and $\rho(0)\approx|j,-j\rangle\langle j,-j|$, then
$\lim_{T\rightarrow 0}\theta_{U}$ $=\arg\langle j,-j|\mathcal{P}\mathrm{e}^{\frac{1}{1+|\zeta|^{2}}{\oint}_{\gamma}J_{z}\left(\bar{\zeta}\mathrm{d}\zeta-\zeta\mathrm{d}\bar{\zeta}\right)}|j,-j\rangle$
$=\mathrm{i}\frac{j}{1+|\zeta|^{2}}\oint_{\gamma}\left(\bar{\zeta}\mathrm{d}\zeta-\zeta\mathrm{d}\bar{\zeta}\right),$ (70)
where $J_{z}|j,-j\rangle=-j|j,-j\rangle$ has been applied. This result precisely reduces to the Berry phase of the ground state $|j,-j\rangle$ [56]. More specifically, when $j=\frac{3}{2}$ and $\theta=\frac{\pi}{2}$, $|\zeta|=1$ and thus $\theta_{U}=\mathrm{i}\frac{3}{4}{\oint}_{\gamma}(-2\mathrm{i}\tan^{2}\frac{\theta}{2})\mathrm{d}\phi=3\pi\equiv\pi\mod 2\pi$, which confirms the result shown in Fig. 1 as $T\rightarrow 0$.