### v4

# Mixed-state geometric phases of coherent and squeezed spin states

Abstract
Two mixed-state geometric phases, known as the Uhlmann phase and interferometric geometric phase (IGP), of spin coherent states (CSSs) and spin squeezed states (SSSs) are analyzed. Exact solutions and numerical results of selected examples are presented. For the $j=3/2$ CSS, the Uhlmann phase exhibits finite-temperature topological phase transitions characterized by abrupt jumps. The IGP for the same state similarly shows discontinuous jumps as the temperature varies. In the case of the $j=1$ one-axis SSS, both Uhlmann phase and IGP display discrete finite-temperature jumps. By contrast, the $j=1$ two-axis SSS shows no such transitions because the Uhlmann phase and IGP both vary smoothly with temperature. We also briefly discuss potential realizations and simulations related to these phenomena in spin systems.
I Introduction
Geometric phases have been a key concept in topological properties of quantum systems [1, 2]. The Berry phase arises when a quantum system undergoes adiabatic cyclic evolution in the context of pure quantum states and has been fundamental in studying topological insulators and superconductors, where the Berry curvature characterizes the underlying topological features [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. However, quantum systems are often in mixed states due to finite temperatures or non-equilibrium conditions, which call for generalizations of the Berry phase to mixed quantum states. A mathematically rigorous approach is the Uhlmann phase [17, 18, 19, 20] for mixed states. Similar to the Berry phase associated with the holonomy of the Berry bundle, the Uhlmann phase is obtained from parallel transport in a cyclic process and reflects the holonomy of the Uhlmann bundle. Moreover, the Uhlmann-Berry correspondence shows that the Uhlmann phase in the low-temperature limit generally agrees with the Berry phase of the ground state [21]. However, the Uhlmann bundle is trivial [22], which limits its applicability to defining other quantized topological indices. Nevertheless, the Uhlmann phase of bosonic and fermionic coherent states have been studied [21], in addition to that of two-level and spin-$j$ systems [23, 24, 25]. With a suitable generalization, the Uhlmann phase of dynamic systems has also been studied [26].
Another approach called the interferometric geometric phase (IGP) for mixed states has been proposed [27] and further analyzed [28, 29, 30, 31, 32, 33], extending the geometric phase concept via an analogy of the optical Mach-Zehnder interferometer [34]. This IGP has been experimentally observed using nuclear magnetic resonance and polarized neutrons [35, 36, 37, 34]. Unlike the Uhlmann phase, the IGP does not require the evolution of the system to be cyclic [27]. In some cases, it can be shown that the IGP is the argument of the weighted sum of the Berry-phase factors of the constituent states [27].
Ref. [38] presents a comparison between the Uhlmann phase and IGP to contrast their differences. The Uhlmann phase, for instance, requires the ancillary system modeling environmental effects to be manipulated according to the change of the system and can exhibit finite-temperature topological phase transitions with quantized jumps in two-level systems. In contrast, the IGP does not exhibit such transitions at finite temperatures for general two-level systems. Recent studies have demonstrated that the IGP remains intact as temperature varies for the Kitaev chain [39], in contrast to the discrete jumps seen in the Uhlmann phase [23]. Nevertheless, quantized jumps of the IGP as temperature increases can occur in three-level or higher-dimensional systems [38]. Interestingly, these generalizations of the mixed-state geometric phases provide further insights into topological quantum phase transitions (TQPTs) and dynamical quantum phase transitions (DQPTs) [40].
Meanwhile, spin systems have served as paradigms in various physics branches, ranging from quantum mechanics [41] to statistical physics [42] to spintronic devices [43]. Two types of spin states have received particular attention: The coherent spin states (CSSs) are minimal uncertainty states of collective angular momentum. The CSS is the building block for the path-integral description of many-body spin systems [44] and serves as a classical approximation of quantum spin systems, thereby establishing the standard quantum limit (SQL) for measurements [45, 46, 47, 48]. However, surpassing the SQL is critical for advancing quantum metrology and quantum information technologies. Spin squeezing [49, 50, 51] allows for a reduction of the variance of one spin component at the expense of increasing that of another orthogonal direction, resulting in the squeezed spin states (SSSs). The SSSs enable precision measurements beyond the SQL and approach the Heisenberg limit with potential applications in, for example, atomic clocks [52, 53]. Significant improvements in precision measurement using spin squeezing have also been made in atomic Bose-Einstein condensates (BECs) [54, 55]. These developments not only highlight the metrological advantage of the SSS but also open possible avenues for generating quantum entanglement in many-body systems, fostering further advancement in quantum information processing [56].
Complementing the progress in surpassing the SQL through spin squeezing, here we set to investigate the geometric phases of the CSS and SSS at finite temperatures and search for topological phases or transitions. The possibility of finding different mixed-state geometric phases according to different evolution conditions already demonstrates rich physics in the underlying topology of finite-temperature systems. While the geometric phase of the ground-state CSS has been analyzed [57, 58], here we explicitly calculate the Uhlmann phase and the IGP for selected CSS and SSS at finite temperatures. We found that for the $j=3/2$ CSS, both Uhlmann phase and IGP demonstrate temperature-dependent topological or geometric phase transitions with quantized jumps. For the $j=1$ one-axis SSS, both phases show similar finite-temperature transitions. In contrast, the $j=1$ two-axis SSS displays a smoothly varying Uhlmann phase with increasing temperature, indicating the absence of topological phase transitions. Likewise, the IGP of that case also evolves continuously with temperature.
The rest of the paper is organized as follows. Sec. II provides a brief, self-contained theoretical background, covering the CSS, SSS, Uhlmann phase, and the adiabatic IGP. Sec. III and Sec. IV present our systematic investigations of the mixed-state geometric phases for selected cases of the CSS and SSS, respectively. Section V discusses possible experimental and theoretical implications of our findings. Finally, Sec. VI concludes our work.
II THEORETICAL BACKGROUND
We begin with a brief overview of the CSS and SSS as well as a short review of the Uhlmann phase and interferometric geometric phase (IGP) of mixed states. Additionally, we will generalize the concept of IGP to allow arbitrary unitary evolution. For simplicity, we set $c=\hbar=k_{\text{B}}=1$ throughout the paper.
II.1 Overview of CSS
The spin operators satisfy the algebra [41]
$[J_{z},J_{+}]=J_{+},\quad[J_{z},J_{-}]=-J_{-},\quad[J_{+},J_{-}]=2J_{z}.$ (1)
The eigenstates of $J^{2}$ and $J_{z}$ are $|jm\rangle,\leavevmode\nobreak\ -j\leq m\leq j$. Moreover, $J_{\pm}|jm\rangle=\sqrt{(j\mp m)(j\pm m+1)}|jm\pm 1\rangle$ and $J_{z}|jm\rangle=m|jm\rangle$. The CSS is defined as [59, 60]
$|\xi\rangle:=\mathrm{e}^{\xi J_{+}-\bar{\xi}J_{-}}|j,-j\rangle,$ (2)
where $\xi\in\mathds{C}$ is a complex number.
Using the Baker-Campbell-Hausdorff disentangling formula [59], the CSS can also be expressed as
$|\xi\rangle$ $=\mathrm{e}^{\zeta J_{+}}\mathrm{e}^{\ln\left(1+|\zeta|^{2}\right)J_{z}}\mathrm{e}^{-\bar{\zeta}J_{-}}|j,-j\rangle$
$=\frac{1}{(1+|\zeta|^{2})^{j}}\mathrm{e}^{\zeta J_{+}}|j,-j\rangle\equiv|\zeta\rangle$ (3)
with $\zeta=\zeta(\xi)=\frac{\xi\tan|\xi|}{|\xi|}$. After defining $D(\xi)\equiv D(\zeta)=\mathrm{e}^{\xi J_{+}-\bar{\xi}J_{-}}$, the CSS corresponds to the ground state of the translated Hamiltonian $\hat{H}(\xi)\equiv\hat{H}(\zeta)=D(\zeta)\hat{H}D^{\dagger}(\zeta)$. Here $\hat{H}=\omega_{0}J_{z}$ , and $\omega_{0}$ is the Larmor frequency. The associated excited states are obtained in a similar manner: $|jm,\xi\rangle=|jm,\zeta\rangle=D(\zeta)|jm\rangle$, $-j\leq m\leq j$. By parameterizing $\xi$ as $\xi=\mathrm{e}^{-\mathrm{i}\phi}\frac{\theta}{2}$, we have $\zeta=\mathrm{e}^{-\mathrm{i}\phi}\tan\frac{\theta}{2}$. If the parameter ranges are chosen as $0\leq\phi\leq 2\pi$, $0\leq\theta\leq\pi$, then $\zeta$ spans the entire complex plane. Consequently, we will use $\zeta$ as the control parameter. The map $(\theta,\phi)\rightarrow\zeta$ projects the unit sphere onto the entire complex plane. Actually, the CSSs are closely related to the eigenstates of the corresponding spin-$j$ system with the Hamiltonian $H=\omega_{0}\hat{\mathbf{B}}\cdot\mathbf{J}$, where $\hat{\mathbf{B}}$ is the unit vector pointing along the direction of an external magnetic field [44].
II.2 Overview of SSS
[…]
$J_{\alpha}=\frac{1}{2}\sum_{l=1}^{N}\sigma_{l\alpha},\quad\alpha=x,y,z,$ (4)
where $\sigma_{l\alpha}$ is the Pauli matrix for the $l$-th object. Here we consider the spin-squeezed states (SSS) generated by the one-axis twisting Hamiltonian [50, 61]
$H_{\mathrm{OAT}}=\eta J_{x}^{2}=\frac{\eta}{4}\sum_{k,l=1}^{N}\sigma_{kx}\sigma_{lx},$ (5)
which is a nonlinear operator with the coupling constant $\eta$ involving all pairwise interactions. This indicates the presence of pairwise correlations in the spin-squeezed states generated by the Hamiltonian. Choosing the initial state as $|j,-j\rangle=|1\rangle^{\otimes N}$, where $|1\rangle=|j=1/2,j_{z}=1/2\rangle$, the SSS at time $t$ is formally written as:
$|\Psi(t)\rangle=\exp\left(-\mathrm{i}\Theta(t)J_{x}^{2}/2\right)|1\rangle^{\otimes N},$ (6)
where $\Theta(t)=2\eta t$.
By defining
$S(\Theta)=\mathrm{e}^{-\mathrm{i}\Theta J_{x}^{2}/2},$ (7)
we can generalize this concept of spin-squeezed states
$|jm,\Theta\rangle=S(\Theta)|jm\rangle.$ (8)
We also introduce
$\tilde{S}(\Theta(t))=\sum_{m}\mathrm{e}^{-\int^{t}_{0}\mathrm{d}t^{\prime}\langle m(t^{\prime})|\frac{\mathrm{d}}{\mathrm{d}t^{\prime}}|m(t^{\prime})\rangle}|m(t)\rangle\langle m(0)|,$ (9)
where $|m(t)\rangle:=|jm,\Theta\rangle$, so that $|m(0)\rangle=|jm\rangle$. When compared to $S(\Theta)$, $\tilde{S}(\Theta)$ simply incorporates a pure-state geometric phase $\mathrm{i}{\int}^{t}_{0}\mathrm{d}t^{\prime}\langle m(t^{\prime})|\frac{\mathrm{d}}{\mathrm{d}t^{\prime}}|m(t^{\prime})\rangle$ to each eigenstate. It is straightforward to verify that $|jm,\Theta\rangle$ is an eigenstate of the Hamiltonian
$H(\Theta)$ $=$ $S(\Theta)\omega_{0}J_{z}S^{\dagger}(\Theta)$ (10)
$=$ $\tilde{S}(\Theta(t))\omega_{0}J_{z}\tilde{S}^{\dagger}(\Theta(t))$
since the geometric phase factors cancel out for each $|jm\rangle$. This in turn implies that the time-evolved thermal state satisfies $\rho(t)=S(\Theta(t))\rho(0)S^{\dagger}(\Theta(t))=\tilde{S}(\Theta(t))\rho(0)\tilde{S}^{\dagger}(\Theta(t))$. The advantage of introducing $\tilde{S}$ will become clear in the subsequent discussions.
Another spin squeezing method is based on a generalized two-axis counter-twisting Hamiltonian [49, 50], which reads
$H_{\mathrm{TAT}}=\mathrm{i}(zJ^{2}_{+}-\bar{z}J^{2}_{-}).$ (11)
Here $z\in\mathds{C}$ is a complex number which can be parameterized as $z=\mathrm{e}^{-\mathrm{i}\phi}\tan\left(\frac{\theta}{2}\right)$ with the parameter ranges chosen as $0\leq\phi\leq 2\pi$, $0\leq\theta\leq\pi$. This is a quadratic squeezing Hamiltonian and it will degrade to the ordinary two-axis Hamiltonian $\frac{\gamma}{2\mathrm{i}}(J_{+}^{2}-J_{-}^{2})$ if $z$ is a purely real number equal to $-\gamma/2$. The squeezing operator $K(z)$ now is defined as
$K(z)=\mathrm{e}^{zJ^{2}_{+}-\bar{z}J^{2}_{-}},$ (12)
which is somewhat similar with the translation operator $D(\xi)=\mathrm{e}^{\xi J_{+}-\bar{\xi}J_{-}}$ we encountered in the discussion of the CSS. However, there are subtle differences in the quadratic form of angular momentum operators. Similar to the case of one-axis squeezing, the eigenstates $|jm,z\rangle$ can be given from the corresponding Hamiltonian
$H(z)=K(z)\omega_{0}J_{z}K^{\dagger}(z).$ (13)
II.3 Uhlmann phase
A mathematically rigorous approach for generalizing the Berry phase from pure states to mixed states is the Uhlmann phase [17] through purification of density matrices, which can be described by the underlying fiber-bundle structure [62], An operator $W$ is said to purify the density matrix $\rho$ if $\rho=WW^{\dagger}$. Conversely, the purification $W$ can be uniquely decomposed as $W=\sqrt{\rho}\mathcal{U}$ if $\rho$ is full-rank (i.e., rank$(\rho)=N$ for a system with a $N$-dimensional Hilbert space). Here the unitary operator $\mathcal{U}$ is referred to as the phase factor of mixed states, representing a generalization of the U(1) phase factor associated with pure states.